OBSERVATIONS OF N_{2}O_{5} AND NO_{3} IN AMBIENT AIR VIA CONTINUOUS-WAVE CAVITY RING-DOWN SPECTROSCOPY

\underline{WILLIAM R. SIMPSON}, Department of Chemistry and Geophysical Institute, University of Alaska Fairbanks, Fairbanks, AK 99775-6160.

The nitrate radical, NO_{3}, and its related reservoir compound, dinitrogen pentoxide, N_{2}O_{5} are important intermediates in atmospheric chemistry. The nitrate radical reacts reversibly with nitrogen dioxide to form dinitrogen pentoxide, and these three compounds are thought to be equilibrated in the ambient atmosphere. These molecules act as intermediates between NO_{x} (NO and NO_{2}) and nitric acid. The nitrate radical, due to its strong visible absorption bands in the red, has been measured in the ambient atmosphere via long-path spectroscopy. Recently, cavity ring-down spectroscopy has been used to measure NO_{3} in-situ, also affording detection of N_{2}O_{5} via thermal dissociation to NO_{3} at 80^{∘}C. We have applied continuous-wave CRDS to detect these compounds in ambient air. This technique will be discussed. We report observations of N_{2}O_{5} in Fairbanks, AK during wintertime. Atmospheric implications of the observations are also discussed.